The estimated coefficients for COVID-19 in British Columbia are shown below where D is a dummy...

  1. The estimated coefficients for COVID-19 in British Columbia are shown below where D is a dummy taking the value of 1 after lockdowns and social distancing around March 20th, and 0 before. What is the growth rate of COVID-19 after the lockdown?

Coefficients

Standard Error

Intercept

2.65924

0.061369

Days

0.202875

0.00567

D

3.264957

0.122193

D*days

-0.1669

0.006542
Answer #1

Let Y be the growth rate of COVID-19.

By the question

After lockdown

=2.65924+0.202875*Days+3.264957-0.1669*Days

Before lockdown

  =2.65924+0.202875*Days

=3.264957- 0.1669*Days